OpenStudy (anonymous):

how do i find a point that is perpendicular to another line if I know what the equation of the first line is and i know what the slope is.

jimthompson5910 (jim_thompson5910):

say the given slope is 2/3 Any line perpendicular to the given line will have a slope of -3/2 One way to remember this is to think "flip the fraction and flip the sign" So.... Start with 2/3 (given slope) flip the fraction: 2/3 ----> 3/2 flip the sign: 3/2 ----> -3/2 Therefore, if the given line has a slope of 2/3, then the slope of any perpendicular line is -3/2

OpenStudy (anonymous):

Im sorry I dont really understand this how do I do the whole problem so that I can do the next 3

jimthompson5910 (jim_thompson5910):

what's the full problem again?

jimthompson5910 (jim_thompson5910):

what's all that you're given?

OpenStudy (anonymous):

If the line h passes through the two points (1, 2) and (3, 6) This is the given information to solve the two problmes, here is the questions What is the slope of the line h passing through the two points? What is the equation of the line h passing through the two points? If line g passes through point (2, 2) and is perpendicular to the line h, then what is the equation of line g ?

jimthompson5910 (jim_thompson5910):

how far did you get here?

jimthompson5910 (jim_thompson5910):

were you able to find the slope of the line h?

OpenStudy (anonymous):

i got that the equation of line h passing through is Y=2x

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the slope is 2 or 2/1

OpenStudy (anonymous):

and the slpe is 2

jimthompson5910 (jim_thompson5910):

so what's the perpendicular slope?

OpenStudy (anonymous):

If line g passes through point (2, 2) and is perpendicular to the line h, then what is the equation of line g ? this is the one that I am having trouble with

jimthompson5910 (jim_thompson5910):

if the slope is 2 or 2/1 then what's the perpendicular slope?

jimthompson5910 (jim_thompson5910):

hint: look at my steps in my very first post

OpenStudy (anonymous):

I really do not like math i am so slow to learn this stuff

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

do you see how in my example (in my first post) where I went from 2/3 to -3/2 did that all make sense?

OpenStudy (anonymous):

I was thinking that Y=-1/2X

jimthompson5910 (jim_thompson5910):

the perpendicular slope is -1/2, that's a good start

jimthompson5910 (jim_thompson5910):

so m = -1/2

jimthompson5910 (jim_thompson5910):

the point line g goes through is (2,2) so this means that x = 2 and y = 2

OpenStudy (anonymous):

but how do I figure this out IM not so good with all these formulas can you dumb this down a bit for me.

jimthompson5910 (jim_thompson5910):

now we turn to y = mx+b and plug in m = -1/2, x = 2 and y = 2 y = mx + b y = (-1/2)x + b ... plug in m = -1/2 2 = (-1/2)*(2)+ b ... plug in x = 2 and y = 2 2 = (-1/2)*(2/1)+ b 2 = (-1*2)/(2*1)+ b 2 = -2/2 + b 2 = -1 + b 2 + 1 = b 3 = b b = 3 Therefore, m = -1/2 and b = 3 So the equation of line g is \[\large y = -\frac{1}{2}x + 3\]

jimthompson5910 (jim_thompson5910):

let me know what you think

OpenStudy (anonymous):

Im still trying to undersdand MX+B i know that those numbers have a funtion and you need to know what those are to move on right?

jimthompson5910 (jim_thompson5910):

m is the slope of the line b is the y-intercept of the line

OpenStudy (anonymous):

that is what my book says

jimthompson5910 (jim_thompson5910):

x and y are variables they can be any number the value of y will depend on what x is

OpenStudy (anonymous):

and you need to do some math to find the slope and I needed to know the other parts to do this problem right?

jimthompson5910 (jim_thompson5910):

the usefulness of y = mx+b is that you can find the equation of any line if you know the slope and any point on the line

OpenStudy (anonymous):

are you a teacher?

jimthompson5910 (jim_thompson5910):

I'm a student learning to become one so yes and no

jimthompson5910 (jim_thompson5910):

and yes, you need to figure out the slope before you can move on

OpenStudy (anonymous):

I think I gave that to you right, which I got 2 for that

OpenStudy (anonymous):

how did you find out y is that one of the four numbers that you use to plot the line?

jimthompson5910 (jim_thompson5910):

what do you mean?

jimthompson5910 (jim_thompson5910):

yes the original slope (ie the slope of h) is 2

OpenStudy (anonymous):

I am trying to understand the whole equation I hate just getting the answers to a problem and not know how to solve it you know, i was wondering what Y is and how do you get the numbers for all the letters in this math problem.

OpenStudy (anonymous):

ive been on google all night looking up how to solve for all these math problems and have been comming up with the right answers but have been struggling to understand them.

jimthompson5910 (jim_thompson5910):

y is a variable and it is not assigned a fixed value like m or b are

jimthompson5910 (jim_thompson5910):

when you say y = 2x, you are saying "for each x you plug in , you double the value to get y" so that means if you plug in x = 1, you get y = 2 (double 1 to get 2) when you plug in x = 13, you get y = 26 (double 13 to get 26)

jimthompson5910 (jim_thompson5910):

this produces a bunch of points, you plot these points, then draw a line through them to graph y = 2x

OpenStudy (anonymous):

My problem is im trying to do these in my head I dont have a calculator or graphing paper :(

jimthompson5910 (jim_thompson5910):

well if you can't do them in your head, use pencil and paper

jimthompson5910 (jim_thompson5910):

you can use google as a calculator ex: type in "2*3 + 10" into google without quotes and it will spit out the result like a calculator

OpenStudy (anonymous):

i was trying to graph the lines to visualize them

jimthompson5910 (jim_thompson5910):

and if you need a program to help you graph, I recommend geogebra http://www.geogebra.org/cms/en/

jimthompson5910 (jim_thompson5910):

to graph the lines, you just need two points so one point on y = 2x is (1,2) another point on y = 2x is (2,4) plot the two points (1,2) and (2,4) and draw a straight line through them this will give you the graph of y = 2x

OpenStudy (anonymous):

here is the next part to the questions that you have helped me with, If line F is parallel to the line H, and line F has Y-intercept 5, what is the equation of line F? can you walk me through all the steps and ill write them down as we go i really want to lean this type of math.

OpenStudy (anonymous):

im guessing it is using the answer to the last question that you helped with.

jimthompson5910 (jim_thompson5910):

what is the slope of line h

jimthompson5910 (jim_thompson5910):

yes it all builds up

OpenStudy (anonymous):

ok let me see what I have for that. I got 2 for that.

jimthompson5910 (jim_thompson5910):

so m = 2

OpenStudy (anonymous):

ok got it.

jimthompson5910 (jim_thompson5910):

the y-intercept is given to be 5, so this means b = 5

jimthompson5910 (jim_thompson5910):

so we go from y = mx+b to y = 2x + 5

jimthompson5910 (jim_thompson5910):

that's the equation of line f

jimthompson5910 (jim_thompson5910):

why is the slope 2? because parallel lines have equal slopes

jimthompson5910 (jim_thompson5910):

slope of h is 2 line f is parallel to h so the slope of line f is also 2

OpenStudy (anonymous):

do you have a aol sceen name i could add you to my list to ask another question some time?

OpenStudy (anonymous):

im still writing down your words for the problem right now.

jimthompson5910 (jim_thompson5910):

I don't use AOL messenger, but I do use yahoo messenger

OpenStudy (anonymous):

oh i dont have that installed now. but you can email me at doclovett@yahoo.com

OpenStudy (anonymous):

So Y and B are the same as 5? is this correct.

jimthompson5910 (jim_thompson5910):

b is 5, y is a variable y is allowed to change

OpenStudy (anonymous):

so y is an unknown

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

it represents any number pretty much at this point the same can be said for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

damn now I have to deal with a triangle and COS and TAN

OpenStudy (anonymous):

thanks for your help. I am trying to figure out angles with Sin A=15/17

jimthompson5910 (jim_thompson5910):

check out this page http://openstudy.com/updates/521ef027e4b0750826e0c91a I think you (zombiebacon) and the user jaidyn206 are taking the same course. I could be wrong.

jimthompson5910 (jim_thompson5910):

so you guys can set up a study session if that is true

OpenStudy (anonymous):

yeah she is in my class im talking to her in aim she is ahead of me in the class on different questions.

jimthompson5910 (jim_thompson5910):

I gotcha

OpenStudy (anonymous):

what does sin A=15/17 mean

OpenStudy (anonymous):

i just wasted a few min figuring out degrees in angles for nothing didn't I?

jimthompson5910 (jim_thompson5910):

if you have a right triangle that looks like this 

jimthompson5910 (jim_thompson5910):

then the sine of the angle A is equal to the opposite side (15) over the hypotenuse 17 ie, sin(A) = 15/17

jimthompson5910 (jim_thompson5910):

no I don't think you wasted your time

OpenStudy (anonymous):

how do I draw a picture to show you

jimthompson5910 (jim_thompson5910):

use the draw button below the text box

jimthompson5910 (jim_thompson5910):

that works